OpenStudy (anonymous):

Find the missing length. Round to the nearest tenth, if necessary. a = 9, c = 17, b = ?

OpenStudy (anonymous):



OpenStudy (anonymous):

You know about the Pythagoras Theorem, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

oh ya!

OpenStudy (anonymous):

Fill in those blanks with the numbers you have.

OpenStudy (anonymous):

\[9^2 + b^2 = 17^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What does your equation look like now?

OpenStudy (anonymous):

\[81 + b = 289\] is what it should look like.

OpenStudy (anonymous):

ok mine does

OpenStudy (anonymous):

Now simply 289 - 81 = \[289 - 81 = \sqrt{b}\]

OpenStudy (anonymous):

well it doesn't equal the square root of b... but you are supposed to find the square root of b after you solve for 289 - 81

OpenStudy (anonymous):

Did you get your answer yet?

OpenStudy (anonymous):

The answer to 289 - 81 is 208

OpenStudy (anonymous):

now find the square root of 208 and that will equal the length of side B.

OpenStudy (anonymous):

14.4?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem, Geometry is a pain, but once you get used to it... it gets easier.

OpenStudy (anonymous):

Ya i hope so!